Heavy Rain Corporation just paid a dividend of $3.42 per share, and the firm is expected to experience constant growth of 5.04% over the foreseeable future. The common stock is currently selling for $63.85 per share. What is Heavy Rain’s cost of retained earnings using the Gordon Model (DDM) approach?
Answer #1

Last Dividend, D0 = $3.42
Growth Rate, g = 5.04%
Current Price, P0 = $63.85

Expected Dividend, D1 = D0 * (1 + g)
Expected Dividend, D1 = $3.42 * 1.0504
Expected Dividend, D1 = $3.592368

Cost of Retained Earnings = D1 / P0 + g
Cost of Retained Earnings = $3.592368 / $63.85 + 0.0504
Cost of Retained Earnings = 0.0563 + 0.0504
Cost of Retained Earnings = 0.1067 or 10.67%